# Repeated Padding as Data Augmentation for Sequential Recommendation

### v1

@@ 3. Experiments
In this section, we first add RepPad to several well-known sequential models to verify whether it can effectively improve recommendation performance. Afterwards, we compare RepPad with representative data augmentation methods, including heuristic and training-required methods. Finally, we explore and analyze the different variants of RepPad.
3.1. Experimental Setting
3.1.1. Datasets.
@@ 3.1. Experimental Setting
  • Toys, Beauty, Sports, and Home: Four datasets obtained from Amazon (McAuley et al., 2015) and correspond to the ”Toys and Games”, ”Beauty”, ”Sports and Outdoors”, and ”Home” categories respectively.
@@
3.1.2. Baselines.
To provide a comprehensive evaluation of RepPad, the baselines consist of three main categories:
@@
Table 4. Performance improvement for on five short-sequence datasets. The ‘w/ RP’ represents adding our method to the corresponding model. The best performance of the two is Bolded. All improvements are significant (p¡0.05 with paired t-tests).
@@
Table 5. Performance improvement for on long-sequence datasets. The ‘w/ RP’ represents adding our method to the corresponding model. The best performance of the two is Bolded. All improvements are significant (p¡0.05 with paired t-tests).
@@
Table 6. Performance comparison between RepPad and heuristic data augmentation methods. The best performance of is bolded and the second best is underlined. Hit@5 and NDCG@5 are omitted due to space constraints.
@@
Home Hit@10 0.0057 0.0064 0.0072 0.0087 0.0078 0.0094 0.0108 Home Hit@10 0.0155 0.0156 0.0166 0.0208 0.0167 0.0173 0.0204
Hit@20 0.0106 0.0120 0.0141 0.0151 0.0149 0.0159 0.0182 Hit@20 0.0244 0.0247 0.0270 0.0336 0.0275 0.0276 0.0311
NDCG@10 0.0027 0.0030 0.0034 0.0042 0.0037 0.0044 0.0054 NDCG@10 0.0083 0.0086 0.0089 0.0109 0.0092 0.0099 0.0110
NDCG@20 0.0039 0.0045 0.0051 0.0059 0.0054 0.0060 0.0073 NDCG@20 0.0101 0.0105 0.0111 0.0141 0.0113 0.0122 0.0137
Second. The second category is the widely used stochastic data augmentation methods, which do not require training but include several hyper-parameters that need to be manually tuned. Random (Ran): Augment each sequence by randomly selecting items from the whole item set $\mathcal{V}$ (Liu et al., 2023). Slide Window (SW): It adopts slide windows to intercept multiple new subsequences from the original sequence. Random-seq (Ran-S): Select items from the original sequence randomly as the augmented items (Liu et al., 2023). CMR: Three augmentation operators (Crop, Mask, and Reorder) used in (Xie et al., 2022). CMRSI: Five augmentation operators (Crop, Mask, Reorder, Substitute, and Insert) used in (Liu et al., 2021a).
Third. The third category is representative of training-required data augmentation models, which usually include trainable parameters and hyper-parameters that need to be manually tuned. ASReP (Liu et al., 2021b) employs a reversely pre-trained transformer to generate pseudo-prior items for short sequences. Then, fine-tune the pre-trained transformer to predict the next item. DiffuASR (Liu et al., 2023) adapts the diffusion model to the item sequence generation. Besides, two guide strategies are designed to control the model to generate the items more corresponding to the raw sequence. CL4SRec (Xie et al., 2022) utilizes contrastive learning to extract self-supervised signals from raw sequential data. We do not include method L2Aug (Wang et al., 2022) since it is designed to augment casual user sequences. CASR (Wang et al., 2021) does not provide open-source codes for reliable reproduction.
3.1.3. Evaluation Metrics.
3.1.4. Implementation Details.
For all baselines, we adopt the implementation provided by the authors. We set the embedding size to 64 and the batch size to 256. We carefully set and tune all other hyper-parameters of each method as reported and suggested in the original papers. We use the Adam (Kingma and Ba, 2014) optimizer with a learning rate 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$. For all methods, we adopt early stopping on the validation set if the performance does not improve for 20 epochs and report results on the test set. We conduct three runs and report the average results for all methods.
@@ 3.2. Overall Improvements
(1) As a whole, the RNN- and CNN-based models perform comparably. GRU4Rec is the best performer in both categories, which verifies the effectiveness of the gated recurrent units module in the sequential recommendation. The Transformer-based models perform better overall, especially SASRec and STOSA. This shows the abilities of the Transformer architecture in modeling user sequence patterns. FMLP-Rec performs slightly less well, which we believe may be related to the fact that the pure MLP architecture cannot accurately capture the fine-grained relationship between each user interaction and historical behavior as the Transformer does.
(2) As demonstrated in Table 4, after adding our RepPad to the original model, the performance of all models was significantly improved. The average recommendation performance improvement is up to 60.3% on GRU4Rec, 24.3% on SASRec, and 22.8% on STOSA. In particular, Our method gives GRU4Rec a boost of up to 110.00% Hit@5 and 122.22% for NDCG@5 on the Home dataset. The above experimental results validate the effectiveness of our method on short-sequence datasets. It also shows that RepPad can be applied to various types of sequential models. Considering that in real-world scenarios, the vast majority of user history records are short sequences (Liu et al., 2021b; Jiang et al., 2021), our approach has great potential for application.
(3) From Table 5, we can observe that RepPad can only achieve less improvement on some models for long-sequence datasets and even cause performance degradation on some methods. These results verify our conjecture in Section 2.5.2. Even though we have set the maximum sequence length to 200 for long-sequence datasets, most of the user sequences still do not have enough space left for repeated padding. At the same time, repeated padding may interfere with the model’s ability to capture long-term user preferences and impair recommendation performance. We also note that the Transformer-based model still outperforms other types of models overall, further demonstrating the power of its architecture.
Figure 2. The result of the ablation study.
3.3. Compare with Other Methods
In this subsection, we compare RepPad with different data augmentation methods. For the heuristic methods, we choose the two representative models, SASRec and GRU4Rec, as the backbone network. For methods that require training, we chose the most versatile SASRec since different methods have different restrictions on the type of backbone network. We present the results in Table 6 and 7.
From Table 6, we observe that RepPad outperforms existing heuristic data augmentation methods in most cases. CMRSI performs the best among the baselines, thanks to the five well-designed data augmentation operators that help mine more preference information while mitigating data sparsity. Ran-S also achieves competitive performance in some examples due to its augmentation using items in the original sequence. We also observed that in some cases, the model performance degraded after using data augmentation methods, which may be related to the fact that these methods introduce too much noise that destroys the integrity of the original sequence. RepPad utilizes the free data space and uses the original sequences for padding, which improves model performance while retaining the user’s complete preference knowledge. From Table 7, RepPad still achieves satisfactory results in the face of data augmentation methods that require training. Among the baseline methods, DiffuASR and CL4SRec leverage the diffusion model and sequence-level augmentation with contrastive learning to achieve competitive performance in some cases. Surprisingly, RepPad achieves the best or second-best performance in all cases without including any parameters and training process.
@@ 3.4. Ablation Study
In this section, we conduct an ablation study to explore the different variants of RepPad. Based on the padding times $m$ in Table 1 and whether or not to add the delimiter 0, We compare the recommendation performance of the following variants: 1) RP-Fix: Performing repeated padding with a fixed number of times, we report the best results for times in range $\{1,2,3\}$. 2) RP-Max: Performing repeated padding until the remaining space is insufficient (i.e. maximum padding times). 3) RP(0, Max): When performing repeated padding each time, randomly select between 0 and the maximum number of times. 4) RP(1, Max): When performing repeated padding each time, randomly select between 1 and the maximum number of times. Fig. 2 presents the variants and their performance. We report the performance of each of the four variants with and without the addition of the delimiter 0.
For performance between different variants, in most results, the performance of RP(0, Max) and RP(1, Max) can come with considerable improvements. RP-Fix and RP-Max bring limited performance improvement compared to RP(0, Max) and RP(1, Max). For PR-Fix, RepPad is executed with a fixed number of times, so it does not fully utilize the idle input space. For RP-Max, we observe that it not only does not bring improvement in some cases but also leads to a significant decrease in model performance. If every sequence performs RepPad to the maximum number of times, it may make the representation learned by the model unbalanced. Because the repeated padding time tends to be large for short sequences, multiple iterations of short behavioral sequences may interfere with the model’s ability for long-sequence and long-term preferences modeling, which in turn impairs recommendation performance. For RP(0, Max) and RP(1, Max), they avoid the imbalance problem caused by RP-Max while fully utilizing the free input space. Meanwhile, RP(1, Max) performs better than RP(0, Max), indicating that executing RepPad at least once is essential.
@@
4. Why RepPad Works
In this section, we provide analyses of why RepPad is effective.
4.1. Training Efficiency
In Fig. 3, we illustrate the loss convergence of different methods during training. We also compare the epoch time consumption of different methods in Table 8. From the figures and tables, we can observe that RepPad’s losses converge slightly faster than the original model. For CMRSI and CL4SRec, the heuristic data augmentation introduces more training samples along with more noise, resulting in more epochs for the model to converge. The contrastive learning task of CL4SRec also imposes additional computational overhead and convergence burden. Also, RepPad only adds a small amount of epoch elapsed time compared to the original model. While CMSRI and CL4SRec add extra training samples and auxiliary tasks, resulting in a substantial increase in training elapsed time. Comparing RP-Fix and RP(1, Max), the former can only be performed repeated padding a fixed and smaller number of times. At the same time, the latter randomly chooses between one time and the maximum number of times, which slightly increases the elapsed time but improves the speed of loss convergence and model performance.
4.2. Gradient Stability
Research has shown that more stable or smooth gradients are more helpful for model training (Li et al., 2019; Reddi et al., 2019; Xu et al., 2019), thus improving model robustness and performance. In Fig. 4, we illustrate the average gradient of item embeddings throughout the training process with and without our RepPad. We show the frequency of the gradient values as histograms. We can observe that the average gradient values are more centralized with the addition of our method. While the gradient is more likely to have extremely large or small values during the original model training, RepPad mitigates this phenomenon, making the gradient more stable during the training process. The RepPad increases the number of training samples input into the model, especially for short interaction sequences. The model can derive more stable gradient estimates from more samples, thus accurately modeling user preferences.
@@ 4.2. Gradient Stability
Method Time Method Time
@@ 4.3. Summary and Insights
Based on the analysis above, we believe that RepPad utilizes the idle input space to provide more samples for model training and improve the training efficiency. RepPad can make the gradient update of the model more stable without increasing the total number of training sequences. Unlike existing methods, our method augments short sequences to a length comparable to that of long sequences in a nearly lossless manner. It enables the model to focus fairly on short and long sequences, giving more attention to long-tail items. In this way, the model can better learn the representation of cold-start users and items and thus give more accurate recommendations. In addition, we speculate that RepPad acts as some kind of regularization during model training.
@@ 5.1. Sequential Recommendation
Sequential recommendation (SR) is an emerging topic in the field of recommender systems (Qu et al., 2022). An early solution is to treat the item sequence as a Markov Chain (Rendle et al., 2010; Shani et al., 2005; He and McAuley, 2016), where the next item to predict is closely related to the latest few interactions. The limitation lies in the inability to learn the dependency in a relatively large time step. Hence, a better solution based on recurrent neural networks (RNNs) (Schuster and Paliwal, 1997) emerges since it can capture long- and short-term preferences. A pioneering work GRU4Rec (Hidasi et al., 2015a) introduced RNN into the field of recommender systems. It trained a Gated Recurrent Unit (GRU) architecture to model the evolution of user interests. Following this idea, CmnRec (Qu et al., 2022) proposed to divide proximal information units into chunks, whereby the number of memory operations can be significantly reduced. Another line of research is based on convolutional neural networks (CNNs). The representative work, Caser (Tang and Wang, 2018), embeds a sequence of recent items in the time and latent spaces, then learns sequential patterns using convolutional filters. NextItNet (Yuan et al., 2019) combines masked filters with 1D dilated convolutions to model the dependencies in interaction sequences. More recently, transformer-based approaches have attracted much attention because of their convenience for parallelization and carefully designed self-attention architecture (Vaswani et al., 2017; Devlin et al., 2018). SASRec (Kang and McAuley, 2018) employed unidirectional Transformers to fulfill the next-item prediction task in the sequential recommendation. Some works further optimized the Transformer’s architecture to make it more suitable for recommendation scenarios, such as reducing computational complexity (Fan et al., 2021) and stochastic self-attention (Fan et al., 2022).
@@ 5.2. Data Augmentation for SR
Data augmentation has been established as a standard technique to train more robust models in many domains. In recommender systems, it has also been widely used and proved to be an effective method. Slide windows is an early work (Tan et al., 2016b), which modified the original click sequence by splitting it into many sub-sequences or randomly discarding a certain number of items. Inspired by counterfactual thinking, CASR (Wang et al., 2021) revised the sequence of user behaviors by substituting some previously purchased items with other unknown items. Some researchers proposed to leverage bi-directional transformers (Liu et al., 2021b; Jiang et al., 2021) and diffusion models (Liu et al., 2023) for sequence data augmentation. Besides, L2Aug (Wang et al., 2022) categorizes users into core and casual users, using historical data from the former to augment the interaction sequence of the latter.
With the rise of self-supervised techniques (e.g., contrastive learning), many researchers used it as an auxiliary task to further improve model performance (Qin et al., 2023; Xie et al., 2022). For example, CL4SRec (Xie et al., 2020) came up with three data augmentation operators (Crop, Reorder, Mask) for pre-training recommendation tasks equipped with a contrastive learning framework. Based on CL4SRec, CoSeRec (Liu et al., 2021a) integrated item similarity information for data augmentation (Insert, Substitute) with a contrastive learning objective, aiming to maximize the agreement of augmented sequences. To further improve the quality of the augmented sequence, some works performed augmentation concerning time and item category information (Dang et al., 2023a; Tian et al., 2023). In addition to sequence-level augmentation, augments and contrasts at the attribute level (Zhou et al., 2020), embedding level (Qiu et al., 2022), and model level (Hao et al., 2023) has also been explored. Unlike existing methods, we are the first to explore the use of idle padding space for data augmentation. In terms of simplicity and applicability, our method does not contain any parameters or training processes and can be plugged into most existing sequential models.

### v2

@@ 3. Experiments
3.1. Experiment Setup
3.2. Datasets
@@
  • Toys, Beauty, Sports, and Home: Four datasets obtained from Amazon (McAuley et al., 2015) and correspond to the ”Toys and Games”, ”Beauty”, ”Sports and Outdoors”, and ”Home” categories, respectively.
@@
3.3. Baselines
In order to conduct a comprehensive evaluation of the effectiveness and superiority of RepPad, the baselines consist of three main categories:
@@ 3.3. Baselines
Second. The second category is the widely used stochastic data augmentation methods, which do not require training but include several hyper-parameters that need to be manually tuned.
  • Random (Ran) (Liu et al., 2023): This method augments each sequence by randomly selecting items from the whole item set $\mathcal{V}$.
  • Slide Window (SW) (Tan et al., 2016b): It adopts slide windows to intercept multiple new subsequences from the original sequence.
  • Random-seq (Ran-S) (Liu et al., 2023): It selects items from the original sequence randomly as the augmentation items.
  • CMR (Xie et al., 2022): This work proposes three sequence data augmentation operators with contrastive learning. In this category, we only use the three operators including Crop, Mask, and Reorder.
  • CMRSI (Liu et al., 2021a): Based on CMR, this work proposes two additional operators with contrastive learning. We adopt five operators including Crop, Mask, Reorder, Substitute, and Insert.
Third. The third category is representative of training-required data augmentation models, which usually include trainable parameters and hyper-parameters that need to be manually tuned.
  • ASReP (Liu et al., 2021b): This method employs a reversely pre-trained transformer to generate pseudo-prior items for short sequences. Then, fine-tune the pre-trained transformer to predict the next item.
  • DiffuASR (Liu et al., 2023): It adopts the diffusion model for sequence generation. Besides, two guide strategies are designed to control the model to generate the items corresponding to the raw data.
  • CL4SRec (Xie et al., 2022): This method leverages random data augmentation and utilizes contrastive learning to extract self-supervised signals from the original data.
We do not include method L2Aug (Wang et al., 2022) and CASR (Wang et al., 2021) since they do not provide open-source codes for reliable reproduction.
3.4. Evaluation Metrics
3.5. Implementation Details
For all baselines, we adopt the implementation provided by the authors. We set the embedding size to 64 and the batch size to 256. To ensure fair comparisons, we carefully set and tune all other hyper-parameters of each method as reported and suggested in the original papers. We use the Adam (Kingma and Ba, 2014) optimizer with the learning rate 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$. For all methods, we adopt early stopping on the validation set if the performance does not improve for 20 epochs and report results on the test set.
Table 4. Performance improvement on the five short-sequence datasets. The ‘w/ RP’ represents adding our method. All improvements are statistically significant, as determined by a paired t-test with $p\leq 0.05$.
@@ 3.5. Implementation Details
Table 5. Performance improvement on the long-sequence dataset. The ‘w/ RP’ represents adding our method.
@@
Table 6. Performance comparison between RepPad and heuristic data augmentation methods. All improvements are statistically significant, as determined by a paired t-test with the second best result in each case ($p\leq 0.05$).
@@
Home Hit@10 0.0057 0.0064 0.0072 0.0087 0.0078 0.0094 0.0108 Home H@10 0.0155 0.0156 0.0166 0.0208 0.0167 0.0173 0.0204
H@20 0.0106 0.0120 0.0141 0.0151 0.0149 0.0159 0.0182 H@20 0.0244 0.0247 0.0270 0.0336 0.0275 0.0276 0.0311
N@10 0.0027 0.0030 0.0034 0.0042 0.0037 0.0044 0.0054 N@10 0.0083 0.0086 0.0089 0.0110 0.0092 0.0099 0.0109
N@20 0.0039 0.0045 0.0051 0.0059 0.0054 0.0060 0.0073 N@20 0.0101 0.0105 0.0111 0.0141 0.0113 0.0122 0.0137
@@ 3.6. Overall Improvements
(1) As a whole, the RNN- and CNN-based models perform comparably. NARM is usually the best performer, which verifies the effectiveness of combining attention mechanisms and RNN architectures. The Transformer-based models perform better overall, especially SASRec and STOSA. This shows the ability of the Transformer architecture to model user sequence patterns. FMLP-Rec performs slightly less well, which we believe may be related to the fact that the pure MLP architecture cannot accurately capture the fine-grained relationship between each user interaction and historical behavior as the Transformer does.
(2) As demonstrated in Table 4, after adding our RepPad to the original model, the performance of all models is significantly improved. The average recommendation performance improvement is up to 60.3% on GRU4Rec, 24.3% on SASRec, and 22.8% on STOSA. In particular, Our method gives GRU4Rec a boost of up to 110.00% Hit@5 and 122.22% for NDCG@5 on the Home dataset. The above experimental results validate the effectiveness of our method on short-sequence datasets. It also shows that RepPad can be applied to various types of sequential models. Considering that in real-world scenarios, the vast majority of user history records are short sequences (Liu et al., 2021b; Jiang et al., 2021), our approach has great potential for application.
(3) From Table 5, we can observe that RepPad can only achieve less improvement on some models for the long-sequence dataset and even cause performance degradation on some methods. These results verify our conjecture in Section 2.5.2. Even though we have set the maximum sequence length to 200 for long-sequence datasets, most user sequences still do not have enough space left for repeated padding. At the same time, repeated padding may interfere with the model’s ability to capture long-term user preferences and impair recommendation performance. We also note that the Transformer-based model still outperforms other types of models overall, further demonstrating the power of its architecture.
@@
Figure 2. The result of the ablation study.
3.7. Compare with Other Methods
In this subsection, we compare RepPad with different data augmentation methods. For the heuristic methods, we choose the two representative models, SASRec and GRU4Rec, as the backbone network. For methods that require training, we chose the most versatile SASRec since different methods have different restrictions on the type of backbone network.
From Table 6, we observe that RepPad outperforms existing heuristic data augmentation methods in most cases. CMRSI performs the best among the baselines, thanks to the five well-designed data augmentation operators that help discover more preference information while mitigating data sparsity. Ran-S also achieves competitive performance in some examples due to its augmentation using items in the original sequence. We also observe that in some cases, the model performance degraded after using data augmentation methods, which may be related to the fact that these methods introduce too much noise that destroys the integrity of the original sequence. RepPad utilizes the free data space and uses the original sequences for padding, which improves model performance while retaining the user’s complete preference knowledge. From Table 7, RepPad still achieves satisfactory results in the face of data augmentation methods that require training. Among the baseline methods, DiffuASR and CL4SRec leverage the diffusion model and sequence-level augmentation with contrastive learning to achieve competitive performance in some cases. Surprisingly, RepPad achieves the best or second-best performance in all cases without including any parameters and training process.
@@ 3.8. Ablation Study
In this section, we conduct an ablation study to explore the different variants of RepPad. Based on the padding times $m$ in Table 1 and whether or not to add the delimiter 0, we compare the performance of the following variants: 1) RP-Fix: Performing repeated padding with a fixed number of times, we report the best results for times in range $\{1,2,3\}$. 2) RP-Max: Performing repeated padding until the remaining space is insufficient (i.e., maximum padding times). 3) RP(0, Max): When performing repeated padding each time, randomly select between 0 and the maximum number of times. 4) RP(1, Max): When performing repeated padding each time, randomly select between 1 and the maximum number of times. Figure 2 presents the results. We report the performance of each of the four variants with and without adding the delimiter 0.
For performance between different variants, in most results, the performance of RP(0, Max) and RP(1, Max) can come with considerable improvements. RP-Fix and RP-Max bring limited performance improvement compared to RP(0, Max) and RP(1, Max). For PR-Fix, RepPad is executed with a fixed number of times, so it does not utilize the idle input space. For RP-Max, we observe that it not only does not bring improvement in some cases but also leads to a significant decrease in model performance. If every sequence performs RepPad to the maximum number of times, it may make the representation learned by the model unbalanced. Because the repeated padding time tends to be large for short sequences, multiple iterations of short behavioral sequences may interfere with the model’s ability for long-sequence and long-term preferences modeling, which in turn impairs recommendation performance. For RP(0, Max) and RP(1, Max), they avoid the imbalance problem caused by RP-Max while utilizing the free input space. Meanwhile, RP(1, Max) performs better than RP(0, Max), indicating that executing RepPad at least once is essential.
@@
Method Epoch Time Method Epoch Time
@@
4. Why RepPad Works
4.1. Loss Convergence and Training Efficiency
In Figure 3, we illustrate the loss convergence of different methods during training. We also compare the epoch time consumption of different methods in Table 8. From the figures and tables, we can observe that RepPad’s losses converge slightly faster than the original model. For CMRSI, the heuristic data augmentation introduces more different training sequences along with more noise, resulting in more epochs for the model to converge. The contrastive learning task of CL4SRec also imposes additional computational overhead and convergence burden. Also, RepPad only adds a small amount of epoch elapsed time compared to the original model. While CMSRI and CL4SRec add extra training sequences and auxiliary tasks, resulting in a substantial increase in training elapsed time. Comparing RP-Fix and RP(1, Max), the former can only be performed repeated padding a fixed and smaller number of times. At the same time, the latter randomly chooses between one time and the maximum number of times, which slightly increases the elapsed time but improves the speed of loss convergence and model performance.
4.2. Gradient Stability
Research has shown that more stable or smooth gradients are more helpful for model training (Li et al., 2019; Reddi et al., 2019; Xu et al., 2019), thus improving model robustness and performance. In Figure 4, we illustrate the average gradient of item embeddings throughout the training process with and without our RepPad. We show the frequency of the gradient values as histograms. We can observe that the average gradient values are more centralized with the addition of our method. While the gradient is more likely to have large or small values during the original model training, RepPad mitigates this phenomenon, making the gradient more stable during the training process. Smooth updating of the gradient allows training to converge stably and improves the model’s recommendation performance
@@ 4.3. Summary and Insights
RepPad can make the gradient update of the model more stable without increasing the total number of training sequences. Unlike existing methods, our method augments short sequences to a length comparable to that of long sequences in a nearly lossless manner. Existing methods inevitably generate noise or compromise the integrity of the original data when perform augmentation. From another perspective, RepPad enables the model to focus fairly on short and long sequences, giving more attention to long-tail items and users. In this way, the model can better learn the representation of cold-start users and items. We also speculate that RepPad acts as some kind of regularization, which mitigates the negative effects of the uneven distribution of original dataset.
@@ 5.1. Sequential Recommendation
Sequential recommendation (SR) is an emerging topic in the field of recommender systems (Qu et al., 2022). An early solution is to treat the item sequence as a Markov Chain (Rendle et al., 2010; Shani et al., 2005; He and McAuley, 2016), where the next item to predict is closely related to the latest few interactions. The limitation lies in the inability to learn the dependency in a relatively large time step. Hence, a better solution based on recurrent neural networks (RNNs) (Schuster and Paliwal, 1997) emerges since it can capture long- and short-term preferences. A pioneering work GRU4Rec (Hidasi et al., 2015a) introduced RNN into the field of recommender systems. It trained a Gated Recurrent Unit (GRU) architecture to model the evolution of user interests. Following this idea, CmnRec (Qu et al., 2022) proposed to divide proximal information units into chunks, whereby the number of memory operations can be significantly reduced. More recently, transformer-based approaches have attracted much attention because of their convenience for parallelization and carefully designed self-attention architecture (Vaswani et al., 2017; Devlin et al., 2018). SASRec (Kang and McAuley, 2018) employed unidirectional Transformers to fulfill the next-item prediction task in the sequential recommendation. Some works further optimized the Transformer’s architecture to make it more suitable for recommendation scenarios, such as reducing computational complexity (Fan et al., 2021) and stochastic self-attention (Fan et al., 2022). Furthermore, FMLP-Rec (Zhou et al., 2022) proposes an all-MLP architecture with learnable filters to enhance recommendation performance.
@@ 5.2. Data Augmentation for SR
Data augmentation has been established as a standard technique to train more robust models in many domains. In recommender systems, Slide Windows is an early work (Tan et al., 2016b), which modified the original click sequence by splitting it into many sub-sequences or randomly discarding a certain number of items. Inspired by counterfactual thinking, CASR (Wang et al., 2021) revised the sequence of user behaviors by substituting some previously purchased items with other unknown items. Some researchers proposed to leverage bi-directional transformers (Liu et al., 2021b; Jiang et al., 2021) and diffusion models (Liu et al., 2023) for sequence data augmentation. With the rise of self-supervised techniques (e.g., contrastive learning), many researchers used it as an auxiliary task to further improve model performance (Qin et al., 2023; Xie et al., 2022). For example, CL4SRec (Xie et al., 2020) developed three data augmentation operators for pre-training recommendation tasks equipped with a contrastive learning framework. Based on CL4SRec, CoSeRec (Liu et al., 2021a) integrated item similarity information for data augmentation with a contrastive learning objective, aiming to maximize the agreement of augmented sequences. Unlike existing methods, we are the first to explore the use of idle padding space for data augmentation. In terms of simplicity and applicability, our method does not contain any parameters or training processes and can be plugged into most existing sequential models.